OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!! Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 9 3x + 4y = 26

OpenStudy (dan815):

a ball costs 2 pieces of cat poo and a bat costs 1 piece of dog poo or 9 dollars all together, but a 3 pieces of cat poo and 4 pieces of dog poo give you 26 dollars in return if you see to a shop, so how much is a piece of dog poo and cat poo worth? part 2 x=(9-y)/2 3((9-y)/2)+4y=26 solve for y :(27-3y)/2+4y=26 5y=52-27 y=5 x=(9-5)/2=2 a piece of dog poo = 5 dollars a piece of cat poo = 2 dollars